OpenStudy (anonymous):

factor z^3+125 im stuck on this

OpenStudy (anonymous):

\[a^3 + b^3 = (a+b)(a^2 +b^2 - ab)\]

OpenStudy (anonymous):

You can write 125 as \(5^3\)..

OpenStudy (anonymous):

Using that formula now calculate \[z^3 + 5^3 = ??\] a = z and b = 5.

OpenStudy (anonymous):

@jen89 , Go ahead.. What happened??

OpenStudy (anonymous):

ummm im stuck well lost

OpenStudy (anonymous):

Lost?? Where??

OpenStudy (anonymous):

See, I have written the formula in my very first post.. There just replace a by z and by 5.. and then tell me what you got?

OpenStudy (anonymous):

z^3+5^2(z+5)(z^2+5^2-z-5)

OpenStudy (anonymous):

did i do that part right

OpenStudy (anonymous):

\[z^3 + 5^3 = (z+5)(z^2 + 5^2 - z \cdot 5)\]

OpenStudy (anonymous):

Now solve this further..

OpenStudy (anonymous):

What is \(5^2\) ??

OpenStudy (anonymous):

25

OpenStudy (anonymous):

So finally you will get: \[z^3 + 5^3 = \color{green}{(z+5)(z^2 + 5^2 - 5z)}\]

OpenStudy (anonymous):

ok

hero (hero):

@waterineyes, try teaching how to factor quadratic trinomials.

OpenStudy (anonymous):

I did not get you @hero..

hero (hero):

Example of quadratic trinomial: \[15x^2 + 19x + 6\]

OpenStudy (anonymous):

To whom I am teaching @hero? You??

hero (hero):

@jen89

OpenStudy (anonymous):

See, here comparing with ax^2 + bx + c a = 15, b = 19 and c = 6 To first find a*c = 15*6 = 90 Now I have to find the factors of 90 such that by adding them I will get 19 ie b.. Example: 2*45 = 90 but by adding 2 + 45, I will not get 19.. So first find two factors of 90 such that first factor + second Factor = 19.. Can you do this??

OpenStudy (anonymous):

Okay, first tell me the possible factors of 90.. Example 15*6 = 90 2*45 = 90 3*30 = 90 Like wise..

OpenStudy (anonymous):

ok 10 and 9 5 and 18

hero (hero):

See that's what I mean

hero (hero):

With the the other method, you have to guess more

OpenStudy (anonymous):

You are telling this to whom?

hero (hero):

To @jen89

OpenStudy (anonymous):

Okay.. Have you come across quadratic polynomial earlier??

OpenStudy (anonymous):

You both

hero (hero):

We've done several together that I went over with her.

hero (hero):

I teach the factor by grouping method as you already know.

hero (hero):

mn = 90 m + n = 19

hero (hero):

Using that method, it's is obvious the two numbers are 10 and 9

OpenStudy (anonymous):

her??

hero (hero):

Then from there, you can proceed to split the middle term

OpenStudy (anonymous):

Is he her??

hero (hero):

Yes....HER

OpenStudy (anonymous):

Ha ha ha.. My mistake..

hero (hero):

Anyway...

hero (hero):

From there, you can split the middle term since 19x = 10x + 9x 15x^2 + 10x + 9x + 6

hero (hero):

Then factor by grouping: 5x(3x + 2) + 3(3x + 2) (3x + 2)(5x + 3)

hero (hero):

If you use any other method, you'll have to guess more

OpenStudy (anonymous):

ok thank you im trying realy hard to get this down pack

OpenStudy (anonymous):

Now can I move from this post??

hero (hero):

What do you mean "move from this post"?

OpenStudy (anonymous):

I mean I have to go. I am in cafe using Openstudy. I have to pay more charges if I don't move.. (

hero (hero):

You don't need my permission to move bro.

OpenStudy (anonymous):

Okay brother.. :) A good enemies can be best friends.. It is a fact.. :) Sorry if I have hurt you ever..

hero (hero):

Trust me, you never hurt me.

OpenStudy (anonymous):

im sorry for all of this

OpenStudy (anonymous):

ok here another one z^3+27 so far i got (z+3) (z^2+3^2-3^2)

hero (hero):

Almost but not quite. The problem is, if you're taking an online test, they will just say it is wrong.

hero (hero):

Also, the way @waterineyes wrote the rule is not the way it should be written

hero (hero):

\[(a^3 + b^3) = (a + b)(a^2 - ab + b^2)\]

OpenStudy (anonymous):

27 is b right

hero (hero):

No, 3 is b

hero (hero):

and 3 cubed is 27

OpenStudy (anonymous):

ok i see how u got the 3 now

OpenStudy (anonymous):

ok i got (z^3+3^3) = (z+3) (a^2-z3+3^2)

hero (hero):

You put a instead of z

hero (hero):

And you should write -3z not -z3

hero (hero):

Well, at least you were able to get this far.

OpenStudy (anonymous):

(z^3+9)=(z+3)(z^2-3z+6)

hero (hero):

Now you're just making it worse. You had everything right except for the a the first time. This time you just made it worse.

hero (hero):

3^2 is not 6 and 3^3 is not 9

OpenStudy (anonymous):

ok i see what i did hold on

OpenStudy (anonymous):

ok 3^2 is 9 3^3 is 27

OpenStudy (anonymous):

(z^3+27)=(z+3)(z^2-3z+9)

OpenStudy (anonymous):

(z^3+27)=(z^3-3z+27) did i do that right

OpenStudy (anonymous):

ok i think i messed that up is it (z^3+27) (z^3-3z^2+3z^2+27)

hero (hero):

This is right: (z^3+27)=(z + 3)(z^2 -3z + 9) You second and third guessing yourself (not good)

hero (hero):

Here's the thing. You need to learn how to multiply polynomials so that you can check your result

hero (hero):

If I multiply (z + 3)(z^2 - 3z + 9) I should be able to get back to z^3 + 27

hero (hero):

That's the only way to truly know if it is right

hero (hero):

So let's do it and see: (z + 3)(z^2 - 3z + 9) = z^2(z + 3) - 3z (z + 3) + 9(z + 3) = z^3 + 3z^2 - 3z^2 - 9z + 9z + 27 = z^3 + 27 Yep, I got back to z^3 + 27, so it must be correct.

hero (hero):

When doing math, the key to being correct is being able to "check" your work. That's the key to mastering it. If you don't know how to check your work, you will never be good at math.

OpenStudy (anonymous):

ok im trying to learn the steps you have been teaching me. i will keep working on this

hero (hero):

Just don't confuse yourself in the process.

OpenStudy (anonymous):

im trying not to

OpenStudy (anonymous):

i feel bad for asking help all the time

hero (hero):

Don't feel bad. Hundreds of students on this site have asked way more questions than you have. But you are actually trying to learn which I appreciate. It's okay to ask questions.

OpenStudy (anonymous):

ok thanks :) by the way i tried one by my self and i want to show you the steps i did to see if i did it right

OpenStudy (anonymous):

X^2+18x+81 mn=81 m+n=18 9 and 9 x(x+9x)+9 (9x+81) (x+9) (x+9)

OpenStudy (anonymous):

i think i checked my work right but im not shore i get the same as i started with

OpenStudy (anonymous):

x^2+9x+9x+81 x^2+18x+81

hero (hero):

You have the right end result, but your intermediary steps are flawed.